OpenStudy (anonymous):

need help asap! WILL MEDAL 2 A cell phone company offers two different monthly text-messaging plans as shown in the table below. For what number of text messages will both plans cost the same for one month? messaging plan\monthly fee\fee per message Plan A $35.00 $0.10 Plan B $18.00 $0.30 A. 55 messages B. 85 messages C. 110 messages D. 140 messages

OpenStudy (anonymous):

Plan A: $35, $0.10 Plan B: $18, $0.30 Let the number of messages be m. $35 + 0.1m = $18 + 0.3m $35 - $18 = 0.3t - 0.1t $17 = 0.2t What is 17 divided by 0.2?

OpenStudy (anonymous):

i get 85

OpenStudy (anonymous):

That's the answer. 85 messages. (:

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem!